A research group conducted an extensive survey of 3146 wage and salaried workers on issues ranging from relationships with their bosses to household chores. The data were gathered through hour-long telephone interviews with a nationally representative sample. In response to the question, "What does success mean to you?" 1540 responded, "Personal satisfaction from doing a good job." Let p be the population proportion of all wage and salaried workers who would respond the same way to the stated question. Find a 90% confidence interval for p. (Round your answers to three decimal places.)

lower limit    
upper limit    
Answer #1

Solution :

Given that,

n = 3146

x = 1540

Point estimate = sample proportion = = x / n = 1540 / 3146 = 0.490

1 - = 1 - 0.490 = 0.510

At 90% confidence level

= 1 - 90%

=1 - 0.90 =0.10

/2 = 0.05

Z/2 = Z0.05 = 1.645

Margin of error = E = Z / 2 * (( * (1 - )) / n)

= 1.645 (((0.490 * 0.510) / 3146)

= 0.015

A 90% confidence interval for population proportion p is ,

± E

= 0.490 ± 0.015

= ( 0.475, 0.505 )

lower limit = 0.475

upper limit = 0.505